The variables x and y vary inversely. Use the given values to write an equation that relates x and y.
x = 15, y = 5

Answer by stanbon(57307)

The variables x and y vary inversely.
y = k/x
---------------------
Use the given values to write an equation that relates x and y.
x = 15, y = 5
---
Solve for "k":
5 = k/15
k = 75
----
Equation:
y = 75/x
==============
Cheers,
Stan H.
==============